OpenStudy (anonymous):

How do i solve for x? 8^(x-2) = 1/16

OpenStudy (anonymous):

write it out as exponentials: (2^3)^(x-2) = 2^(-4) 2^(3x-6) = 2^(-4) 3x-6 = -4 can you solve from here?

OpenStudy (anonymous):

yep, u a life saver bro.

OpenStudy (anonymous):

thx. :) i like to save lives.